OpenStudy (anonymous):

can someone help me with the steps of this so I can understand it better? If f(x) varies directly with x and f(x) = 2 when x = 10, then what is the value of f(x) when x = 40? (Points : 4) 8 45 200 320

jimthompson5910 (jim_thompson5910):

"f(x) varies directly with x" which means f(x) = k*x where k is some constant

jimthompson5910 (jim_thompson5910):

we are told that " f(x) = 2 when x = 10", so plug these two values in and then solve for k f(x) = k*x 10 = k*2 10/2 = k 5 = k k = 5

jimthompson5910 (jim_thompson5910):

That means our equation goes from f(x) = k*x to f(x) = 5x

jimthompson5910 (jim_thompson5910):

Do you see how to finish up?

OpenStudy (anonymous):

ok then would I divide f(x) to x to get the k?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

once you figure out k, you plug in x = 40 into f(x) = 5x to get your final answer

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

sure thing